If I give my brother $5$ dollars, then we will have the same amount of money. If instead he gives me $12$ dollars, then I'll have four times as much money as he will have. How much money does my brother currently have (in dollars)?
Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3816)


Answer: $23.33

Explanation

x = amount of money you start with
y = amount of money your brother starts with

Scenarion 1) You give your brother $5
Scenarion 2) Your brother gives you $12
Your amountYour brother's amount
Scenario 1x-5y+5
Scenario 2x+12y-12

At the conclusion of scenario 1, the two of you have the same amount of money.
your amount = his amount
x-5 = y+5
y = x-5-5
y = x-10
We'll use this equation later in a substitution step.

At the conclusion of scenario 2, you have four times as much money compared to your brother.
your amount = 4*(his amount)
x+12 = 4*(y-12)
x+12 = 4*(y-12)
x+12 = 4*(x-10-12) .... substitution
x+12 = 4(x-22)
x+12 = 4x-88
x-4x = -88-12
-3x = -100
x = -100/(-3)
x = 100/3
x = 33.33 dollars approximately
y = x-10 = 33.33 - 10 = 23.33 dollars is the amount your brother starts with.

Check:
Your amountYour brother's amount
Scenario 133.33-5 = 28.3323.33+5 = 28.33
Scenario 233.33+12 = 45.3323.33-12 = 11.33

45.33/11.33 = 4.000882612533 approximately
We don't land on 4 exactly, but we get close enough.

Answer by ikleyn(52778)

.
If I give my brother $5$ dollars, then we will have the same amount of money. If instead he gives me
$12$ dollars, then I'll have four times as much money as he will have. How much money does my brother
currently have (in dollars)?
~~~~~~~~~~~~~~~~~~~~~~~~


As worded,  printed,  posted and presented,  this problem is mathematically  FATALLY  DEFECTIVE,
since it has no a precise solution in integer number of cents.

In such problems,  their meaning  ALLWAYS  assumes a precise solution;  approximate solutions
with non-integer number of cents are not appropriate,  are not accepted and can not be accepted.